OpenStudy (anonymous):

Macaurin series of 10^x, please

OpenStudy (anonymous):

10*10*10*10*10*10*10*10*10*10

OpenStudy (anonymous):

let f(x)=10^x \[f \prime \left( x \right)=10^x \log 10\] \[f \prime \prime \left( x \right)=10^x \left( \log 10 \right)^2\] \[f \prime \prime \prime \left( x \right)=10^x \left( \log10 \right)^3\] ................................................................. .................................................................. ................................................................... \[f^n \left( x \right)=10^x \left( \log 10 \right)^{n}\] .................................................................... ...................................................................... \[f \left( 0 \right)=10^0=1,f \prime \left( 0 \right)=10^0 (\log 10)=\log10\] \[f \prime \prime \left( 0 \right)=10^0 \left( \log 10 \right)^2\] \[f \prime \prime \prime \left( 0 \right)=10^0\left( \log 10 \right)^3\] ............................................................................................ \[f^n \left( 0 \right)=10^0\left( \log10 \right)^n\] Maclaurin series for f(x) \[f(x)=f(0)+x f \prime \left( 0 \right)+\frac{ x^2 }{ 2! }f \prime \prime \left( 0 \right)+\frac{ x^3 }{ 3! }f \prime \prime \prime \left( 0 \right) +...+\frac{ x^n }{ n! }f^n \left( 0 \right) +...\] put the values and complete it.

OpenStudy (anonymous):

Alternatively, you can get it from the power series for \(e^x\): \[10^x=e^{\ln 10^x}=e^{x\ln10}\] \[e^x=\sum_{n=0}^\infty \frac{x^n}{n!}~~\implies~~10^x=e^{x\ln10}=\sum_{n=0}^\infty \frac{(x\ln10)^n}{n!}\]

TheSmartOne (thesmartone):

Welcome to OS!!